OpenStudy (unklerhaukus):

\[\left|\frac1{h-1}\right|<1\]

OpenStudy (calculator):

\[(\frac{1}{h-1})^2<1^2 \\ \frac{1}{h^2-2h+1}<1 \\ h^2-2h+1>1 \\ h^2-2h>0 \\ h(h-2)>0\]  h<0 h>2

OpenStudy (calculator):

How do you think?

OpenStudy (anonymous):

\[\left|\frac{1}{h-1}\right|<1\]\[\frac{1}{|h-1|}<1\]\[|h-1|>1\]this gives us Either \[h-1>1\]Or\[h-1<-1\]So, h >2 or h <0 \[\[h \epsilon (- \infty,0)U(2,\infty)\]

OpenStudy (praxer):

\[-1<\ \frac{ 1 }{ h-1 } <1\] i suppose the equation proceeds from this step.

OpenStudy (praxer):

because we know when modulus of any equation is less then a real number then the equation can be either less then the positive value of the real number or greater then the negative value of the real number

OpenStudy (unklerhaukus):

Thanks @Calculator, @akitav, @praxer . I see where i was going wrong now .

OpenStudy (calculator):

Differnt approaches :)